What is the ph of a 1.15M solution of methanoic acid? Ka= 1.8*10^4
1
Expert's answer
2020-10-01T07:52:34-0400

The chemical equation:

_________HCOOH + H2O <=> HCOO- + H3O+

ICE table: 1.15________________0______0

___________-x________________ +x_____ +x

__________1.15-x______________ x ______x

"Ka = \frac{[HCOO-][H3O+]}{[HCOOH]} = \frac{x^2}{1.15 – x} = 1.8\times10^-4"

Check for negligibility:

"\frac{1.15}{1.8*10^{-4}} = 6388 > 400"

x<<1.15

Then

"\frac{x^2}{1.15} = 1.8\times10^{-4}"

"x^2 = 1.15\times1.8\times10^{-4}"

x = 0.0143

[H3O+] = 0.0143 mol/L

pH = -log[H3O+] = -log0.0143 = 1.84

Answer: 1.84